OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

@chosenmatt

OpenStudy (alexandervonhumboldt2):

@freckles

OpenStudy (freckles):

well didn't we already talk about the set {0,1} under subtraction? and we didn't select that choice when your question was asking for which operations the set is closed.. anyways let's think about the last set since it is a finite set {0,1,2} is 1-2 in {0,1,2}?

OpenStudy (anonymous):

yup

OpenStudy (freckles):

1-2 is -1 right?

OpenStudy (freckles):

is -1 a member of the set {0,1,2}?

OpenStudy (anonymous):

no

OpenStudy (freckles):

so since 1-2 is no in {0,1,2} then the set {0,1,2} is not closed under subtraction... now rational numbers rational numbers look like a/b or c/d where a,b,c,d are integers (and b and d not equal to 0) \[\frac{a}{b}-\frac{c}{d}\] when you subtract these two fractions do you still get a rational number?

OpenStudy (freckles):

try combining the fractions

OpenStudy (anonymous):

yes you still get a rational #

OpenStudy (freckles):

ok then that would make the rationals closed under subtraction --- now odd natural numbers... odd natural numbers={1,3,5,7,9,11,13,...} try to see if you can think of two numbers that you can subtract from this set and that result is not in the set {1,3,5,7,9,11,13,...}

OpenStudy (anonymous):

13 and 1

OpenStudy (anonymous):

13-1 =12

OpenStudy (freckles):

sounds great :) also 2k+1 and 2m+1 are odd numbers where k and m are integers (2k+1)-(2m+1) 2k+1-2m-1 2k+2m+1-1 2k+2m+0 2k+2m 2(k+m) and an even number is not going to be in any set of odd numbers

OpenStudy (freckles):

so you only had one choice to go with here

OpenStudy (anonymous):

rational, right?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

thx

OpenStudy (freckles):

\[\frac{a}{b}-\frac{c}{d}=\frac{ad-cb}{bd} \in \mathbb{Q}\] subce ad-cb is an integer and bd is also in integer and a rational number is a integer/integer (where the bottom doesn't equal 0 of course)